OpenStudy (anonymous):

One number is greater than half of a another number by 15. What are the 2 numbers if their sum is 48? PLEASEHELP.

OpenStudy (anonymous):

make it into a system of equations

OpenStudy (anonymous):

Can you give me the equation, so i can solve for it?

Directrix (directrix):

Let's try this. Let x be the "another" number Then, the "other" number is (1/2) * x + 15 So, x + (1/2) * x + 15 = 48 1.5x + 15 + 48 ---> @Beatriceroces Solve for x Then, you will need to substitute the value of x you get into (1/2) * x + 15 to get the other number.

Directrix (directrix):

Post what you get for x, and we can compare answers, okay?

OpenStudy (anonymous):

@Directrix Okay :)

Directrix (directrix):

@Beatriceroces What did you get for x?

OpenStudy (anonymous):

@Directrix 64.5 ???? I think my answer is wrong .

Directrix (directrix):

>>>64.5 Not correct. Let's see. x + (1/2) * x + 15 = 48 1.5 x = 48 - 15 1.5x = 33 x = 33/1.5 @Beatriceroces --> Divide 33 by 1.5 to get x. Then, you will need to substitute the value of x you get into (1/2) * x + 15 to get the other number.